# Revealing Trends in Datasets from the 2022 ACL and EMNLP Conferences

### v1

@@ 5 Collaboration in Dataset Construction
Google Research, Bar Ilan University, Intel Labs Israel, Salesforce AI Research, Shanghai Mental Health Center, University of Cambridge, AWS AI Labs, Yonsei University, Seoul National University, University of Warwick, Northeastern University, AstraZeneca, King’s College London, The Alan Turing Institute, Institut Polytechnique de Paris, Télécom-Paris, École Polytechnique, Korea University, Adobe Research, AIGEN Sciences, Hong Kong University of Science and Technology, University of Science and Technology of China, Microsoft Azure Cognitive Services Research, Heidelberg University, SRH Hochschule Heidelberg, Beijing University of Posts and Telecommunications, Soochow University, Fudan University, Baidu Inc, University of Trento, Free University of Bozen-Bolzano, Fondazione Bruno Kessler, University of Dayton Research Institute, SIL International, Masakhane, Carnegie Mellon University, Microsoft Semantic Machines, University of Utah, Meta AI, University of Oregon, Adobe Research, UC San Diego, UMass Amherst, DAMO Academy Alibaba Group, Singapore University of Technology and Design, Hong Kong Polytechnic University, KAIST, SoftlyAI Research, IIT Kharagpur, Goldman Sachs Data Science and ML Group, Leibniz University of Hannover, IIT Patna, Bosch Center for Artificial Intelligence, Heidelberg University, Shanghai AI Laboratory, Huawei Noah’s Ark Lab, University of Washington, The University of Hong Kong, JD AI Research, Umeå University, Adlede AB, Zhejiang University, AZFT Joint Lab for Knowledge Engine, UNC Chapel Hill, Allen Institute for AI, Tencent AI Lab, Renmin University of China, Hong Kong Polytechnic University, Chinese Academy of Sciences, University of Chinese Academy of Sciences, Islamic University of Technology, Chinese University of Hong Kong, Tsinghua University, IT University of Copenhagen, LMU Munich, Munich Center for ML, Shanghai AI Lab, University of Melbourne, RMIT University, MBZUAI, Data Friendly Space, ToggleCorp Solutions, Johannes Kepler University Linz, Stony Brook University, University of Arizona, Stanford University, VinAI Research, Hanoi University of Science and Technology, University of Oxford, University of Texas at Austin, UC Santa Barbara, Southeast University, Nanjing Origin Information Technology Company, IIT Bombay, Uniphore, HuggingFace, Yale University, University of Southern California, Kakao Corp, Soongsil University, Seoul National University, Virginia Tech, Jamia Millia, University of Sydney, University of New South Wales, University of Notre Dame
@@ 7 The Rise of Multimodality
These datasets contain multiple languages (see Table 9). MSCTD (Liang et al., 2022) in English and Chinese; Filighera et al. (2022) in English and German; DiS-ReX (Bhartiya et al., 2022) in English, French, Spanish and German; Maronikolakis et al. (2022) in Brazilian Portuguese, German, Hindi, Swahili and English; XFUND (Xu et al., 2022a) in 7 languages Chinese, Japanese, Spanish, French, Italian, German and Portuguese; Crossmodal-3600 (Thapliyal et al., 2022) in 36 languages; MT-GenEval Currey et al. (2022) in Arabic, French, German, Hindi, Italian, Portuguese, Russian and Spanish; IndicNLG Benchmark Kumar et al. (2022) in 11 Indic languages; EUR-Lex-Sum (Aumiller et al., 2022) in 24 official European languages; ClidSum (Wang et al., 2022b) in English, German and Chinese; BLOOM (Leong et al., 2022) in 363 languages; MEE (Pouran Ben Veyseh et al., 2022) in 8 languages English, Spanish, Portuguese, Polish, Turkish, Hindi, Korean, and Japanese; HumSet (Fekih et al., 2022) in English, Spanish, French;  Lai et al. (2022) in 5 languages English, Danish, Spanish, Turkish, and Urdu; TyDiP (Srinivasan and Choi, 2022) in 9 languages Hindi, Korean, Spanish, Tamil, French, Vietnamese, Russian, Afrikaans, Hungarian; and ClassBases (Wiriyathammabhum, 2022) in English, Spanish, Portuguese.
@@
ClassBases (Wiriyathammabhum, 2022) 3 English, Spanish, Portuguese
@@ Appendix B Dataset details
 Wiriyathammabhum (2022) ClassBases at the CASE-2022 Multilingual Protest Event Detection Task: Multilingual Protest News Detection and Automatically Replicating Manually Created Event Datasets COVID-related protest event datasets from the New York Times news corpus event detection document classification—XLM-RoBERTa-base, mLUKE-base and XLM-RoBERTa-large AND sentence classification— mLUKE-base AND token classification—XLM-RoBERTa-base Yes (English, Spanish and Portuguese) No

### v2

@@ 5 Collaboration in Dataset Construction
Google Research, Bar Ilan University, Intel Labs Israel, Salesforce AI Research, Shanghai Mental Health Center, University of Cambridge, AWS AI Labs, Yonsei University, Seoul National University, University of Warwick, Northeastern University, AstraZeneca, King’s College London, The Alan Turing Institute, Institut Polytechnique de Paris, Télécom-Paris, École Polytechnique, Korea University, Adobe Research, AIGEN Sciences, Hong Kong University of Science and Technology, University of Science and Technology of China, Microsoft Azure Cognitive Services Research, Heidelberg University, SRH Hochschule Heidelberg, Beijing University of Posts and Telecommunications, Soochow University, Fudan University, Baidu Inc, University of Trento, Free University of Bozen-Bolzano, Fondazione Bruno Kessler, University of Dayton Research Institute, SIL International, Masakhane, Carnegie Mellon University, Microsoft Semantic Machines, University of Utah, Meta AI, University of Oregon, Adobe Research, UC San Diego, UMass Amherst, DAMO Academy Alibaba Group, Singapore University of Technology and Design, Hong Kong Polytechnic University, KAIST, SoftlyAI Research, IIT Kharagpur, Goldman Sachs Data Science and ML Group, Leibniz University of Hannover, IIT Patna, Bosch Center for Artificial Intelligence, Heidelberg University, Shanghai AI Laboratory, Huawei Noah’s Ark Lab, University of Washington, The University of Hong Kong, JD AI Research, Umeå University, Adlede AB, Zhejiang University, AZFT Joint Lab for Knowledge Engine, UNC Chapel Hill, Allen Institute for AI, Tencent AI Lab, Renmin University of China, Hong Kong Polytechnic University, Chinese Academy of Sciences, University of Chinese Academy of Sciences, Islamic University of Technology, Chinese University of Hong Kong, Tsinghua University, IT University of Copenhagen, LMU Munich, Munich Center for ML, Shanghai AI Lab, University of Melbourne, RMIT University, MBZUAI, Data Friendly Space, ToggleCorp Solutions, Johannes Kepler University Linz, Stony Brook University, University of Arizona, Stanford University, VinAI Research, Hanoi University of Science and Technology, University of Oxford, University of Texas at Austin, UC Santa Barbara, Southeast University, Nanjing Origin Information Technology Company, IIT Bombay, Uniphore, HuggingFace, Yale University, University of Southern California, Kakao Corp, Soongsil University, Seoul National University, Virginia Tech, Jamia Millia, University of Sydney, University of New South Wales, University of Notre Dame, University of Groningen
@@ 7 The Rise of Multimodality
These datasets contain multiple languages (see Table 9). MSCTD (Liang et al., 2022) in English and Chinese; Filighera et al. (2022) in English and German; DiS-ReX (Bhartiya et al., 2022) in English, French, Spanish and German; DivEMT (Sarti et al., 2022) in Arabic, Dutch, Italian, Turkish, Ukrainian, and Vietnamese;  Maronikolakis et al. (2022) in Brazilian Portuguese, German, Hindi, Swahili and English; XFUND (Xu et al., 2022a) in 7 languages Chinese, Japanese, Spanish, French, Italian, German and Portuguese; Crossmodal-3600 (Thapliyal et al., 2022) in 36 languages; MT-GenEval Currey et al. (2022) in Arabic, French, German, Hindi, Italian, Portuguese, Russian and Spanish; IndicNLG Benchmark Kumar et al. (2022) in 11 Indic languages; EUR-Lex-Sum (Aumiller et al., 2022) in 24 official European languages; ClidSum (Wang et al., 2022b) in English, German and Chinese; BLOOM (Leong et al., 2022) in 363 languages; MEE (Pouran Ben Veyseh et al., 2022) in 8 languages English, Spanish, Portuguese, Polish, Turkish, Hindi, Korean, and Japanese; HumSet (Fekih et al., 2022) in English, Spanish, French;  Lai et al. (2022) in 5 languages English, Danish, Spanish, Turkish, and Urdu; TyDiP (Srinivasan and Choi, 2022) in 9 languages Hindi, Korean, Spanish, Tamil, French, Vietnamese, Russian, Afrikaans, Hungarian; and ClassBases (Zavarella et al., 2022; Hürriyetoğlu et al., 2021) in English, Spanish, Portuguese.
@@
ClassBases (Zavarella et al., 2022; Hürriyetoğlu et al., 2021) 3 English, Spanish, Portuguese
@@
DivEMT Sarti et al. (2022) 6 Arabic, Dutch, Italian, Turkish, Ukrainian, Vietnamese
@@ Appendix B Dataset details
 Zavarella et al. (2022); Hürriyetoğlu et al. (2021) ClassBases at the CASE-2022 Multilingual Protest Event Detection Task: Multilingual Protest News Detection and Automatically Replicating Manually Created Event Datasets COVID-related protest event datasets from the New York Times news corpus event detection document classification—XLM-RoBERTa-base, mLUKE-base and XLM-RoBERTa-large AND sentence classification— mLUKE-base AND token classification—XLM-RoBERTa-base Yes (English, Spanish and Portuguese) No